# Quantum algorithms for calculating determinant and inverse of matrix and solving linear algebraic systems

### v3

@@ III Inverse matrix
It is remarkable that the minor modification of the algorithm for calculating the determinant of a square $N\times N$ matrix in Sec.II allows to calculate the inverse of a square $(N-1)\times(N-1)$ matrix. Moreover, this modification leads to a minor increase in the depth of the algorithm which becomes $O(N^{2}\log N)$ instead of $O(N\log^{2}N)$ in the algorithm for calculating the determinant.
III.1 Preliminary arguments
@@
The algorithm discussed below allows to invert the $(N-1)\times(N-1)$ matrix $A$,
$A=\left(\begin{array}[]{ccc}a_{11}&\cdots&a_{1(N-1)}\cr a_{21}&\cdots&a_{2(N-1)}\cr\cdots&\cdots&\cdots\cr a_{(N-1)1}&\cdots&a_{(N-1)(N-1)}\end{array}\right),$ (34)
(rather than the matrix $\tilde{A}$ itself) which is embedded into $\tilde{A}$.
@@
where we take into account that $A_{ij}=a_{ij}$, $i,j=1,\dots,N-1$. In turn, this term includes the following term (remove the 1st row and $j$th column from the determinant (43); the $j$th column contains the elements $1$, $a_{1j}$, $a_{2j}$, …, $a_{(N-1)j}$):
$-\frac{(-1)^{i+j}q}{\sqrt{N}}\left|\begin{array}[]{cccccc}a_{11}&\cdots&a_{1(j-1)}&a_{1(j+1)}&\cdots&a_{1(N-1)}\cr\cdots&\cdots&\cdots&\cdots&\cdots&\cdots\cr a_{(i-1)1}&\cdots&a_{(i-1)(j-1)}&a_{(i-1)(j+1)}&\cdots&a_{(i-1)(N-1)}\cr a_{(i+1)1}&\cdots&a_{(i+1)(j-1)}&a_{(i+1)(j+1)}&\cdots&a_{(i+1)(N-1)}\cr\cdots&\cdots&\cdots\cr a_{(N-1)1}&\cdots&a_{(N-1)(j-1)}&a_{(N-1)(j+1)}&\cdots&a_{(N-1)(N-1)}\end{array}\right|=$ (50)
$-\frac{q}{\sqrt{N}}(-1)^{i+j}M_{ij}=-\frac{\det(A)q}{\sqrt{N}}A^{-1}_{ji}.$
Thus, each element $A^{-1}_{ji}$ of the inverse matrix is included into $\det(\tilde{A})$ up to a factor which is the same for all elements. All what remains is to extract all the terms composing the element $A^{-1}_{ji}$ from those composing $\det(\tilde{A})$ and encode them into the proper state of some quantum subsystem. We show that this can be simply realized.
@@ III.2 Basic algorithm
The circuit for the algorithm is presented in Fig.3.
Figure 3: Structure of a quantum circuit for matrix inversion. Ancillae are green on line.
According to (7), the pure state $|\Phi_{0}\rangle$ of the whole system is given in (9) with normalization (8) and constraints (23) for the elements in the first row and first column of the matrix $\tilde{A}$, see (29). From this state, we select only those terms that compose the determinant of $\tilde{A}$ and therefore contain the terms that appear in the resulting inverse matrix $A^{-1}$:
$|\Phi_{0}\rangle=\sum_{{k_{0},\dots,k_{N-1}=0}\atop{k_{0}\neq\dots\neq k_{N-1}}}^{N-1}a_{0k_{0}}\dots a_{(N-1)k_{N-1}}|k_{0}\rangle_{S_{0}}\cdots|k_{N-1}\rangle_{S_{N-1}}+|g_{1}\rangle_{S}$ (51)
while the garbage $|g_{1}\rangle_{S}$ will be removed at the last step of the algorithm after ancilla measurement. However, not all terms in the first part of $|\Phi_{0}\rangle$ are needed in our algorithm. We have to keep only those of them that include the 1st power of the factor $q$, (see constraints (23), i.e., exactly one state $|0\rangle_{S_{j}}$, $11\leq j\leq N-$ 1, can appear in the product $|k_{1}\rangle_{S_{1}}\dots|k_{N-1}\rangle_{S_{N-1}}$). We also have to associate each of the probability amplitudes in (51) with the appropriate element $A^{-1}_{ij}$ of the inverse matrix. To do so, we introduce two $n$-qubit subsystems $R$ and $C$ whose basis states serve to enumerate, respectively, rows and columns of the inverse matrix $A^{-1}$.
@@
This operator conditionally applies $\sigma^{(x)}_{R^{(j)}C^{(i)}}$ which is the product of $\sigma^{(x)}$ (acting on the subsystems $R$ and $C$) such that $\sigma^{(x)}_{R^{(j)}C^{(i)}}|0\rangle_{R}|0\rangle_{C}=|j\rangle_{R}|i\rangle_{C}$. The depth of $W_{SRC}^{(ij)}$ is $O(n)$. To establish this estimation we use the fact that the controlled operator $W_{SRC}^{(ij)}$ with $n$-qubit controlling register can be represented in terms of Toffoli gates by a circuit of depth $O(n)$ KShV and require that all $\sigma^{(x)}$ operators in $\sigma^{(x)}_{R^{(j)}C^{(i)}}$ can be applied simultaneously. Applying the operator ${\hat{W}}_{SRC}=\prod_{i=1}^{N-1}\prod_{j=1}^{N-1}W_{SRC}^{(ij)}$ (whose depth is $O(N^{2}n)$) to $|\Phi_{0}\rangle|0\rangle_{R}|0\rangle_{C}$ we obtain
$|\Phi_{1}\rangle=\hat{W}_{SRC}|\Phi_{2}\rangle|0\rangle_{R}|0\rangle_{C}=$ (54)
$\frac{1}{\sqrt{N}}\sum_{{i=1}}^{N-1}\sum_{{j=1}}^{N-1}\sum_{{0\neq k_{1}\neq\dots\neq k_{i-1}}\atop{\neq k_{i+1}\neq\dots\neq k_{N-1}\neq j}}a_{1k_{1}}\dots a_{(i-1)k_{i-1}}a_{i0}a_{(i+1)k_{i+1}}\dots a_{(N-1)k_{N-1}}\times$
@@
$=-\frac{q\det(A)}{2^{(\tilde{N}+n)/2}}\sum_{{i=1}}^{N-1}\sum_{{j=1}}^{N-1}A^{-1}_{ji}|j\rangle_{R}|i\rangle_{C}|0\rangle_{S_{0}}|1\rangle_{S_{1}}\otimes|{N-1}\rangle_{S_{N-1}}|0\rangle_{A}+|g_{1}^{(D)}\rangle_{SA}.$
@@
$=-\frac{q\det(A)}{2^{(\tilde{N}+n)/2}}\sum_{{i=1}}^{N-1}\sum_{{j=1}}^{N-1}A^{-1}_{ji}|j\rangle_{R}|i\rangle_{C}|0\rangle_{S}|0\rangle_{A}+|g_{2}^{(D)}\rangle_{SA},$
@@
$|\Phi_{4}\rangle=W^{(2)}_{SAB}|\Phi_{3}\rangle|0\rangle_{B}$ $=$ $\sum_{{i=1}}^{N-1}\sum_{{j=1}}^{N-1}a_{ji}|j\rangle_{R}|i\rangle_{C}|0\rangle_{S}|0\rangle_{A}|1\rangle_{B}+|g\rangle_{SA}|0\rangle_{B},$ (57)
$a_{ji}$ $=$ $-\frac{q\det(A)A^{-1}_{ji}}{2^{(\tilde{N}+n)/2}},$ (58)
@@
$|\Phi_{5}\rangle=|\Psi_{out}\rangle|0\rangle_{S}|0\rangle_{A}.$ (59)
@@
The state $|\Psi_{out}\rangle$ encodes $A^{-1}$ in the normalized form, i.e., $|\Psi_{out}\rangle\langle\Psi_{out}|=1$, while the state $|A^{-1}\rangle=\sum_{{i=1}}^{N-1}\sum_{{j=1}}^{N-1}A^{-1}_{ji}|j\rangle_{R}|i\rangle_{C}$ is not normalized. The depth of discussed algorithm for matrix inversion is determined by the operator $\hat{W}_{SRC}$ and equals $O(N^{2}\log N)$. The circuit includes $N\log N$ qubits of the subsystem $S$, $\tilde{N}=O(N\log N)$ qubits of the ancilla $A$, one qubit of the ancilla $B$ and $2\log N$ qubits of $R$ and $C$, thus the size of this circuit approximately equals to the size of the circuit for calculating the determinant which is $O(N\log N)$.
IV Application: solving systems of linear algebraic equations
Having constructed the inverse matrix $A^{-1}$ encoded into the state of the subsystem $R\cup C$ we obtain a tool for solving a system of linear algebraic equations represented in the following matrix form:
$A\bm{x}=\bm{b}.$ (61)
Its solution reads
$\bm{x}=A^{-1}\bm{b}.$ (62)
Therefore, we have to apply the multiplication algorithm proposed in Ref.ZQKW_arxiv2023 (see also Appendix, Sec.VI.1, for its modified version) to the matrices $A^{-1}$ and $\bm{b}$. The structure of the circuit is shown in Fig.4. More detailed circuit is presented in Fig.5.
Figure 4: Structure of the circuit solving a system of linear algebraic equations. Both blocks share the same ancillae, which are not shown in figure.
Notice that we can use the same ancillae in both blocks of this circuit (ancillae are not shown in Fig.4). Moreover, the subsystem $S$ encoding the matrix $A$ can be used as ancilla in further calculations.
In this case we do not measure ancillae $B$ (unlike Sec.III.2) and thus start with the state $|\Phi_{4}\rangle$ given in (57). The vector $\bm{b}$ is encoded as
@@ IV Application: solving systems of linear algebraic equations
The zero value for $b_{0}$ is prescribed by the encoding of $A^{-1}$ in (60), where the 00-states of $R$ and $C$ are not used. Out of system $S\cup A$ (which form ancilla in Fig.5), we take the one-qubit ancilla $\tilde{B}$ needed in the multiplication algorithm and neglect other qubits. We also apply the operator $X_{B}=\sigma^{(x)}$ to invert the state of $B$. Thus, the state of the whole system reads
$|\Phi^{(L)}_{5}\rangle$ $=$ $X_{B}|\Phi_{4}\rangle|b\rangle|0\rangle_{\tilde{B}}=a_{jl}b_{i}|j\rangle_{R}|l\rangle_{C}|i\rangle_{b}|0\rangle_{B}|0\rangle_{\tilde{B}}+|g_{1}\rangle_{RCb}|1\rangle_{B}|0\rangle_{\tilde{B}}.$ (64)
@@
$W^{(3)}_{CBb\tilde{B}}=P_{CBb}\otimes\sigma^{(x)}_{\tilde{B}}+(I_{CBb}-P_{CBb})\otimes I_{\tilde{B}},\;\;P_{CBb}=|0\rangle_{C}|0\rangle_{B}|0\rangle_{b}\;{{}_{C}}\langle 0|_{B}\langle 0|{{}_{b}}\langle 0|.$ (65)
Applying the operator $W^{(3)}_{CBb\tilde{B}}W^{(2)}_{C}W^{(1)}_{Cb}$ to $|\Phi^{(L)}_{5}\rangle$ (64) we obtain
@@
where $|x\rangle_{R}=\sum_{j=1}^{N-1}a_{j_{1}j}b_{j}|j_{1}\rangle_{R}$ is an unnormalized vector, the normalization $G{=(\sum_{l,j}|a_{lj}b_{j}|^{2})^{1/2}}$ is known from the probability of the above measurement which equals $G^{2}/2^{n}$, where $G^{2}\sim\frac{1}{2^{\tilde{N}}}\sim\frac{1}{2^{O(Nn)}}$ due to the expression (58) for $a_{jl}$. Thus, the probability is $\sim 1/2^{\tilde{N}+n}=2^{-O(Nn)}$. The result is stored in the register $R$. The depth of the whole algorithm is defined by the depth of the subroutine inverting $A$ and thus equals $O(N^{2}n)$. In comparison with the circuit for inverse matrix, the circuit in Fig. 5 includes $\log N$ qubits encoding the column $\bm{b}$ and thus remains essentially $O(N\log N)$.
Figure 5: Circuit for the algorithm solving a system of linear algebraic equations (61). Circuit of matrix inversion (up to the operator $W^{(2)}_{SAB}$) is denoted by a block $A^{-1}$. All input qubits of the matrix-inversion algorithm (system $S$) become ancillae qubits. The output of the matrix-inversion algorithm (systems $R$ and $C$) together with system $b$ encoding the vector $\bm{b}$ become the input of the matrix-multiplication block. The output of the whole algorithm is encoded in the state $|x\rangle_{R}$ of the system $R$. All ancillae qubits are green on line. We omit the superscript $(L)$ at $|\Phi_{k}\rangle$, $k=5,6,7$ and at $|\Psi_{out}\rangle$ for brevity.

### v4

@@ III Inverse matrix
We remark here that the minor modification of the algorithm for calculating the determinant of a square $N\times N$ matrix in Sec.II allows to calculate the inverse of a square $(N-1)\times(N-1)$ matrix. Moreover, this modification leads to a minor increase in the depth of the algorithm which becomes $O(N^{2}\log N)$ instead of $O(N\log^{2}N)$ in the algorithm for calculating the determinant.
III.1 Preliminary consideration
@@
The algorithm discussed below allows to invert the $(N-1)\times(N-1)$ matrix $A$ (rather than the matrix $\tilde{A}$ itself) which is embedded into $\tilde{A}$,
$A=\left(\begin{array}[]{ccc}a_{11}&\cdots&a_{1(N-1)}\cr a_{21}&\cdots&a_{2(N-1)}\cr\cdots&\cdots&\cdots\cr a_{(N-1)1}&\cdots&a_{(N-1)(N-1)}\end{array}\right).$ (34)
@@
In turn, this term includes the following term (remove the 1st row and $j$th column from the determinant (43); the $j$th column contains the elements $1$, $a_{1j}$, $a_{2j}$, …, $a_{(N-1)j}$):
$-\frac{(-1)^{i+j}q}{\sqrt{N}}\left|\begin{array}[]{cccccc}a_{11}&\cdots&a_{1(j-1)}&a_{1(j+1)}&\cdots&a_{1(N-1)}\cr\cdots&\cdots&\cdots&\cdots&\cdots&\cdots\cr a_{(i-1)1}&\cdots&a_{(i-1)(j-1)}&a_{(i-1)(j+1)}&\cdots&a_{(i-1)(N-1)}\cr a_{(i+1)1}&\cdots&a_{(i+1)(j-1)}&a_{(i+1)(j+1)}&\cdots&a_{(i+1)(N-1)}\cr\cdots&\cdots&\cdots\cr a_{(N-1)1}&\cdots&a_{(N-1)(j-1)}&a_{(N-1)(j+1)}&\cdots&a_{(N-1)(N-1)}\end{array}\right|$ (50)
$=-\frac{q}{\sqrt{N}}(-1)^{i+j}M_{ij}=-\frac{\det(A)q}{\sqrt{N}}A^{-1}_{ji}.$
Thus, each element $A^{-1}_{ji}$ of the inverse matrix is included into $\det(\tilde{A})$ upto a factor which is the same for all elements. The final step involves extracting all the terms that constitute the element $A^{-1}_{ji}$ from the terms that make up $\det(\tilde{A})$ and encode them into the appropriate state of some quantum subsystem. We show that this process can be easily realized.
@@ III.2 Basic algorithm
The circuit for the algorithm is presented in Fig.3(a). According to (7), the pure state $|\Phi_{0}\rangle$ of the whole system is given in (9) with normalization (8) and constraints (23) for the elements in the first row and first column of the matrix $\tilde{A}$, see (29). From this state, we select only those terms that compose the determinant of $\tilde{A}$ and therefore contain the terms that appear in the resulting inverse matrix $A^{-1}$, i.e., we consider $|\Phi_{0}\rangle$ in form (10).
Figure 3: (Color online) (a) Quantum circuit designed for matrix inversion (with the ancillae shown in green). (b) Quantum circuit designed for solving a system of linear algebraic equations. Both blocks of the circuit share the same ancillae (not shown in the accompanying figure).
However, not all terms in the first part of $|\Phi_{0}\rangle$ are needed in our algorithm. We have to keep only those of them that include the 1st power of the factor $q$, (see constraints (23)), i.e., exactly one state $|0\rangle_{S_{j}}$, $11\leq j\leq N-$ 1, can appear in the product $|k_{1}\rangle_{S_{1}}\dots|k_{N-1}\rangle_{S_{N-1}}$. We also have to associate each of the probability amplitudes in (10) with the appropriate element $A^{-1}_{ij}$ of the inverse matrix. To do so, we introduce two $n$-qubit subsystems $R$ and $C$ whose basis states serve to enumerate, respectively, rows and columns of the inverse matrix $A^{-1}$.
@@
This operator conditionally applies $\sigma^{(x)}_{R^{(j)}C^{(i)}}$ which is the product of $\sigma^{(x)}$ (acting on the subsystems $R$ and $C$) such that $\sigma^{(x)}_{R^{(j)}C^{(i)}}|0\rangle_{R}|0\rangle_{C}=|j\rangle_{R}|i\rangle_{C}$. The depth of $W_{SRC}^{(ij)}$ is $O(n)$. To establish it we use the fact that the controlled operator $W_{SRC}^{(ij)}$ with $n$-qubit controlling register can be represented in terms of Toffoli gates by a circuit of depth $O(n)$ KShV and require that all $\sigma^{(x)}$ operators in $\sigma^{(x)}_{R^{(j)}C^{(i)}}$ can be applied simultaneously. Applying the operator ${\hat{W}}_{SRC}=\prod_{i=1}^{N-1}\prod_{j=1}^{N-1}W_{SRC}^{(ij)}$ (whose depth is $O(N^{2}n)$) to $|\Phi_{0}\rangle|0\rangle_{R}|0\rangle_{C}$ we obtain
$|\Phi_{1}\rangle=\hat{W}_{SRC}|\Phi_{2}\rangle|0\rangle_{R}|0\rangle_{C}$ (53)
$=\frac{1}{\sqrt{N}}\sum_{{i=1}}^{N-1}\sum_{{j=1}}^{N-1}\sum_{{0\neq k_{1}\neq\dots\neq k_{i-1}}\atop{\neq k_{i+1}\neq\dots\neq k_{N-1}\neq j}}a_{1k_{1}}\dots a_{(i-1)k_{i-1}}a_{i0}a_{(i+1)k_{i+1}}\dots a_{(N-1)k_{N-1}}\times$
@@
$=-\frac{q\det(A)}{2^{(\tilde{N}+n)/2}}\sum_{{i=1}}^{N-1}\sum_{{j=1}}^{N-1}A^{-1}_{ji}|0\rangle_{S_{0}}|1\rangle_{S_{1}}$
$\otimes|{N-1}\rangle_{S_{N-1}}|j\rangle_{R}|i\rangle_{C}|0\rangle_{A}+|g_{1}^{(D)}\rangle_{SA}.$
@@
$=-\frac{q\det(A)}{2^{(\tilde{N}+n)/2}}\sum_{{i=1}}^{N-1}\sum_{{j=1}}^{N-1}A^{-1}_{ji}|0\rangle_{S}|j\rangle_{R}|i\rangle_{C}|0\rangle_{A}+|g_{2}^{(D)}\rangle_{SA},$
@@
$|\Phi_{4}\rangle=W^{(2)}_{SAB}|\Phi_{3}\rangle|0\rangle_{B}$
$=\sum_{{i=1}}^{N-1}\sum_{{j=1}}^{N-1}a_{ji}|0\rangle_{S}|j\rangle_{R}|i\rangle_{C}|0\rangle_{A}|1\rangle_{B}+|g\rangle_{SA}|0\rangle_{B},$ (56)
where
$a_{ji}=-\frac{q\det(A)A^{-1}_{ji}}{2^{(\tilde{N}+n)/2}}.$ (57)
@@
$|\Phi_{5}\rangle=|\Psi_{out}\rangle|0\rangle_{S}|0\rangle_{A},$ (58)
@@
The state $|\Psi_{out}\rangle$ encodes $A^{-1}$ in the normalized form, i.e., $\langle\Psi_{out}|\Psi_{out}\rangle=1$, while the state $|A^{-1}\rangle=\sum_{{i=1}}^{N-1}\sum_{{j=1}}^{N-1}A^{-1}_{ji}|j\rangle_{R}|i\rangle_{C}$ is not normalized. The depth of discussed algorithm for matrix inversion is determined by the operator $\hat{W}_{SRC}$ and equals $O(N^{2}\log N)$. The circuit includes $N\log N$ qubits of the subsystem $S$, $\tilde{N}=O(N\log N)$ qubits of the ancilla $A$, one qubit of the ancilla $B$ and $2\log N$ qubits of $R$ and $C$, thus the space of this circuit is approximated by the space of the circuit for calculating the determinant which is $O(N\log N)$.
We note that this depth is obtained under assumption that the depth of the operator $W^{(1)}_{SA}$ in (15) is $N\log^{2}N$. If this depth is $N^{2}\log^{2}N$, then the depth of the matrix inversion algorithm is also $N^{2}\log^{2}N$.
IV Solving systems of linear algebraic equations
Having constructed the inverse matrix $A^{-1}$ encoded into the state of the subsystem $R\cup C$, we obtain a tool for solving a system of linear algebraic equations represented in the matrix form, $A\bm{x}=\bm{b}$, whose solution reads $\bm{x}=A^{-1}\bm{b}.$
Therefore, we have to apply the multiplication algorithm proposed in Ref.ZQKW_arxiv2023 (see also Appendix, Sec.VI.1, for its modified version) to the matrices $A^{-1}$ and $\bm{b}$. The structure of the circuit is shown in Fig.3(b). More detailed circuit is presented in Fig.4.
Figure 4: (Color online) Circuit for quantum algorithm solving a system of linear algebraic equations $A\bm{x}=\bm{b}$. Circuit of matrix inversion (up to the operator $W^{(2)}_{SAB}$) is denoted by a block $A^{-1}$. All input qubits of the matrix-inversion algorithm (system $S$) become ancillae qubits. The output of the matrix-inversion algorithm (systems $R$ and $C$) together with system $b$ encoding the vector $\bm{b}$ become the input of the matrix-multiplication block. The output of the whole algorithm is encoded in the state $|x\rangle_{R}$ of the system $R$. All ancillae qubits are shown in green color. We omit the superscript $(L)$ in $|\Phi_{k}\rangle$ $(k=5,6,7)$ and $|\Psi_{out}\rangle$ for brevity.
Notice that we can use the same ancillae in both blocks of this circuit (ancillae are not shown in Fig.3(b)). Moreover, the subsystem $S$ encoding the matrix $A$ can be used as ancilla in further calculations.
In this case we do not measure ancillae $B$ (unlike Sec.III.2) and thus start with the state $|\Phi_{4}\rangle$ given in (III.2). The vector $\bm{b}$ is encoded as
@@ IV Solving systems of linear algebraic equations
The zero value for $b_{0}$ is prescribed by the encoding of $A^{-1}$ in (59), where the 00-states of $R$ and $C$ are not used. Out of system $S\cup A$ (which forms an ancilla in Fig.4), we take the one-qubit ancilla $\tilde{B}$ needed in the multiplication algorithm and neglect other qubits. We also apply the operator $X_{B}=\sigma^{(x)}$ to invert the state of $B$. Thus, the state of the whole system reads
$|\Phi^{(L)}_{5}\rangle=X_{B}|\Phi_{4}\rangle|b\rangle|0\rangle_{\tilde{B}}$
$=a_{jl}b_{i}|j\rangle_{R}|l\rangle_{C}|i\rangle_{b}|0\rangle_{B}|0\rangle_{\tilde{B}}+|g_{1}\rangle_{RCb}|1\rangle_{B}|0\rangle_{\tilde{B}}.$ (61)
@@
$P_{CBb}$ $=$ $|0\rangle_{C}|0\rangle_{B}|0\rangle_{b}\;{{}_{C}}\langle 0|_{B}\langle 0|{{}_{b}}\langle 0|,$
$W^{(3)}_{CBb\tilde{B}}$ $=$ $P_{CBb}\otimes\sigma^{(x)}_{\tilde{B}}+(I_{CBb}-P_{CBb})\otimes I_{\tilde{B}}.$ (62)
Applying the operator $W^{(3)}_{CBb\tilde{B}}W^{(2)}_{C}W^{(1)}_{Cb}$ to $|\Phi^{(L)}_{5}\rangle$ (IV) we obtain
@@
where $|x\rangle_{R}=\sum_{j=1}^{N-1}a_{j_{1}j}b_{j}|j_{1}\rangle_{R}$ is an unnormalized vector, the normalization $G{=(\sum_{l,j}|a_{lj}b_{j}|^{2})^{1/2}}$ is known from the probability of the above measurement which equals $G^{2}/2^{n}$, where $G^{2}\sim\frac{1}{2^{\tilde{N}}}\sim\frac{1}{2^{O(Nn)}}$ due to the expression (57) for $a_{jl}$. Thus, the probability is $\sim 1/2^{\tilde{N}+n}=2^{-O(Nn)}$. The result is stored in the register $R$. The depth of the whole algorithm is defined by the depth of the subroutine inverting $A$ and thus equals $O(N^{2}n)$. In comparison with the circuit for inverse matrix, the circuit in Fig4 includes $\log N$ qubits encoding the column $\bm{b}$ and thus remains essentially $O(N\log N)$.